# Thin gap approximations for microfluidic device design

### v1

@@ 4 Breakdown of the 2D model
Although the 2D approximation (7a) shows good accord with 3D simulations, there is mild disagreement in flow features such as eddy size. The framework used to derive the approximation naturally lends itself to determining where and how the approximation breaks down. In forming our approximation we have adopted the Hele-Shaw approximation of assuming a parabolic velocity profile across the gap, and no fluid velocity in the depth direction. Based on Fig. 4, we focus particularly on dissecting the features of moderate Reynolds number flows that cause these approximations to break down. We first studied whether the assumption of zero vertical velocity breaks down. Across the entire centrifuge-on-a-chip and for every value of $\mathrm{Re}$ that we simulated, the flow is nearly two-dimensional: the largest value of $|w|_{\infty}/|\mathbf{u}|_{\infty}$ we observe is only 0.0555, obtained in the chamber outlet.
We then investigate whether the assumption of a parabolic velocity breaks down. We define a non-parabolicity quantifier for the $x$-component of the velocity field:
$r:=\frac{|\left\langle u,\overline{Q_{3}}\right\rangle|}{\sqrt{\left\langle u,\overline{Q_{1}}\right\rangle^{2}+\left\langle u,\overline{Q_{3}}\right\rangle^{2}}},\quad\overline{Q_{k}}\equiv Q_{k}/\sqrt{\left\langle Q_{k},Q_{k}\right\rangle}.$ (8)
$r$ measures the relative contribution of the third and first modes to the Gegenbauer polynomials expansion of $u(x,y,z)$. A larger value of $r$ indicates a greater deviation from the parabolic profile. $r$ values remain small ($\lesssim 10^{-2}$) throughout most of the centrifuge-on-a-chip even at $\mathrm{Re}=200$. Moderate values (up to $10^{-1}-1$) occur, in small patches distributed on both the proximal and distal walls of the chamber and more coherently in a jet of fluid entering the centrifuge via the inlet channel (Fig. 5). Inlet jets are common features of moderate to high-Reynolds number flows, and are associated with the inertia of the fluid entering the centrifuge chamber, which doesn’t decelerate to 0 until the jet has penetrate a distance $\approx 25$ into the centrifuge. The presence of the jet violates the Hele-Shaw assumption in two regards: 1. There is a thin boundary layer in the $xy$-plane between the jet and the nearly still fluid in the chamber, and 2. the flow profile is non-parabolic with shear that is closely localized near the $z=\pm 1/2$ walls (Fig. 5b).
Our coaxial jet results (Fig.3 b) show that narrow $x$- or $y$-dimensions alone do not break the thin-gap approximation; the main limitation is its inability to represent non-parabolic profiles. A natural remedy is to add higher-order terms to the asymptotic expansion. Including just one extra term (e.g., $Q_{3}$) markedly improves accuracy, and a two-term expansion reproduces all velocity profiles observed across the jet boundary layer (Fig.5d). While the parabolic form cannot capture non-convex or boundary-layer shapes, the two-term projection resolves both.
5 Discussion
We introduced a modified MWR to derive optimal closure equations, interpreting the Hele–Shaw approximation as the leading term of an orthogonal polynomial expansion for the velocity. This framework reproduces a previously derived approximation Martin et al. (2002); Li et al. (2021) through a simpler, more direct dominant-term analysis, while naturally extending to higher-order corrections.
@@ 5 Discussion
What benefit can be gained from higher-order approximations? First, in many microfluidic thin-gap flows, the vertical velocity remains small, making truly 3D flows rare in thin-gap geometries. Conversely, truly 2D flows occur only in special cases, such as Stokes flow in a channel with constant radius of curvature Lauga et al. (2004). Some applications, however, deliberately induce strong three-dimensionality, e.g. micromixing Stroock et al. (2002). In such cases, using a higher-order leading form for $w$ from our orthogonal polynomials, e.g., $w=w_{1}(x,y)(Q_{4}(z)-Q_{2}(z))$, can improve accuracy, but yields coupled PDEs without a straightforward closure. Second, higher-order terms provide a built-in mechanism for estimating and controlling approximation error. If the estimated error exceeds a chosen tolerance, additional terms can be systematically added to improve accuracy.
@@
While our analysis primarily focused on the fluid velocity equation, the approach is flexible and broadly applicable. It can be readily extended to other physical quantities, such as solute concentration, electric fields, and magnetic fields, arising in applications like reactive solute transport Salmon & Ajdari (2007), electrokinetic flows Lin et al. (2008); Ding (2023), and magnetohydrodynamic microfluidics Qian & Bau (2009). The method also allows for generalization of boundary conditions—for example, incorporating no-stress boundaries to model bubble motion Booth et al. (2025) or flows in microchannels with soft walls Inamdar et al. (2020).
@@
Declaration of interests. The authors report no conflict of interest.

### v2

@@ 4 Higher order corrections
Although the 2D approximation (19a) shows good accord with 3D simulations, there is mild disagreement in flow features such as eddy size. The framework used to derive the approximation naturally lends itself to determining where and how the approximation breaks down. In forming our approximation, we have adopted the Hele-Shaw approximation of assuming a parabolic velocity profile across the gap, and no fluid velocity in the depth direction. Based on Fig. 4, we focus particularly on dissecting the features of moderate Reynolds number flows that cause these approximations to break down. We first studied whether the assumption of zero vertical velocity breaks down. Across the entire centrifuge-on-a-chip and for every value of $\mathrm{Re}$ that we simulated, the flow is nearly two-dimensional: the largest value of $|w|_{\infty}/|\mathbf{u}|_{\infty}$ we observe is only 0.0555, obtained in the chamber outlet.
Next, we examine the validity of the parabolic–velocity assumption. Although a priori estimates for the breakdown of parabolicity can be obtained through linear perturbation analyses of the 3D Navier–Stokes equations in simple geometries, such results remain limited in scope. For example, Plouraboué & Hinch (2002) analyzed Kelvin–Helmholtz instability in a Hele–Shaw cell and showed that the velocity profile of an inhomogeneous fluid may lose its parabolic character at Reynolds numbers as low as $Re\approx 10$. Because the fluid in our study is homogeneous, the mechanism responsible for this breakdown does not directly apply, and the relevant Reynolds-number threshold may be different.
Our interest is instead in an a posteriori assessment tailored to the Reynolds number and geometry of the microfuge. We quantify deviations from parabolicity by comparing the magnitudes of the first two terms in a Gegenbauer-polynomial expansion of the 3D velocity field.
Specifically, we interpret the Hele–Shaw approximation as the leading term of an orthogonal polynomial expansion for the velocity. In this, we go beyond reproducing previously derived approximations (Martin et al. (2002); Li et al. (2021)), generating higher-order corrections on the models for the first time. Including the next order terms, the velocity field and pressure are approximated by
$\mathbf{u}_{H}=\frac{3}{2}\left(1-4z^{2}\right)\left(\mathbf{u}_{0,H}+\left(30z^{2}-\frac{3}{2}\right)\mathbf{u}_{2,H}\right),$ (20)
[…]
where $\mathbf{f}_{H}$ and $f_{z}$ are the body force terms in the horizontal direction and the vertical direction, included here for generality. In the expansion (20), the horizontal velocity component is taken to be even in $z$, which is the most common situation in microfluidic applications. The expression for $\mathbf{u}_{H}$ follows from retaining both $C_{2}^{(-1/2)}(2z)$ and $C_{4}^{(-1/2)}(2z)$, together with appropriate normalization factors. Due to the no-slip boundary condition and the continuity equation, both $w$ and $\partial_{z}w$ vanish on the boundary, which determines the gapwise basis function that multiplies $w_{2}$. Moreover, $n\geq 4$, $C_{n}^{(-\frac{3}{2})}(2z)=\frac{-3}{2n-3}\left(C_{n}^{(-\frac{1}{2})}(2z)-C_{n-2}^{(-\frac{1}{2})}(2z)\right)$ and their derivatives vanish at $z=\pm\frac{1}{2}$, making them a suitable choice of gapwise basis functions for expanding $w$. Accordingly, the vertical momentum equation should be projected onto $C_{n}^{(\frac{5}{2})}(2z)$. The coefficient of $p_{2}$ is justified by examining the residual of the leading-order approximation (6). When $a=6/5$, the $z$-dependent factor becomes
$\frac{3}{2}\left(1-4z^{2}\right)-a=-\frac{1}{5}\left(30z^{2}-\frac{3}{2}\right)=-\frac{1}{5}C_{2}^{(\frac{3}{2})}(2z).$ (21)
This choice of basis for $p_{2}$ to efficiently compensate for the dominant residual generated by the other fields. Furthermore, since $\partial_{z}\left(C_{n}^{(\frac{3}{2})}(z)\right)=3C_{n-1}^{(\frac{5}{2})}(z)$, the derivative $\partial_{z}p$ naturally produces the $C_{n}^{(5/2)}$ family required to balance the error in the vertical momentum equation.
Adopting the expansions in (20) and applying the same weighted residual analysis, the continuity equation yields
$\nabla_{H}\cdot\mathbf{u}_{0,H}=0,\quad\nabla_{H}\cdot\mathbf{u}_{2,H}=\frac{20}{3}w_{2}.$ (22)
The momentum equation in the horizontal direction gives
$\mathrm{Re}\left(\frac{6}{5}\partial_{t}\left(\mathbf{u}_{0,H}-\frac{3}{7}\mathbf{u}_{2,H}\right)+\frac{24}{7}\left(-\mathbf{u}_{0,H}+\frac{54}{11}\mathbf{u}_{2,H}\right)w_{2}\right$. (23)
$\quad\left.+\frac{54}{35}\left(\mathbf{u}_{0,H}-\frac{2}{3}\mathbf{u}_{2,H}\right)\cdot\nabla_{H}\mathbf{u}_{0,H}+\frac{36}{35}\left(-\mathbf{u}_{0,H}+\frac{24}{11}\mathbf{u}_{2,H}\right)\cdot\nabla_{H}\mathbf{u}_{2,H}\right)$
[…]
$\mathrm{Re}\left(\frac{1}{5}\partial_{t}\left(-\mathbf{u}_{0,H}+4\mathbf{u}_{2,H}\right)-\frac{8}{11}\left(\mathbf{u}_{0,H}+\frac{18}{13}\mathbf{u}_{2,H}\right)w_{2}\right$.
$\quad\left.+\frac{2}{5}\left(-\mathbf{u}_{0,H}+\frac{24}{11}\mathbf{u}_{2,H}\right)\cdot\nabla_{H}\mathbf{u}_{0,H}+\frac{48}{55}\left(\mathbf{u}_{0,H}-\frac{9}{26}\mathbf{u}_{2,H}\right)\cdot\nabla_{H}\mathbf{u}_{2,H}\right)$
[…]
The momentum equation in the vertical direction gives
[…]
We first apply the above equation to the unidirectional shear flow problem considered in Section 2.1. We have $u=\frac{3}{2}\left(1-4z^{2}\right)\left(u_{0}+\left(30z^{2}-\frac{3}{2}\right)u_{2}\right)$, where $u_{0}$ and $u_{2}$ solve
[…]
$0=-\frac{18}{35}\Delta_{H}u_{0}+\frac{72}{35}\Delta_{H}u_{2}-\frac{1296}{7}u_{2}.$
If we set $u_{2}=0$ and only keep the first equation, then the system reduces to the closure equation discussed in Section 2.1. By an appropriate change of variables, we can reduce equation (25) to two decoupled equations, leading to the following solution:
$u_{i}(y)=\sum\limits_{j\in\left\{0,2\right\}}a_{0j}a_{ij}\left(1-\text{sech}\left(\frac{L}{2\sqrt{\lambda_{j}}}\right)\cosh\left(\frac{y}{\sqrt{\lambda_{j}}}\right)\right),\;i=0,2,$ (26)
[…]
$0,\displaystyle a_{00}=-\sqrt{\frac{1}{2}+\frac{4\sqrt{\frac{7}{19}}}{5}}\approx-992,\;a_{02}=\sqrt{\frac{1}{2}-\frac{4\sqrt{\frac{7}{19}}}{5}}\approx 120$,
[…]
When $L=1$, the $L^{\infty}$ error of the second-order approximation is $0.0036$, which is roughly one tenth of the error associated with the leading-order approximation described in Section 2.1. The second-order approximation, shown as the orange curve in Figure 2(a), is visually indistinguishable from the exact solution. As illustrated in Figure 2(b,c), the errors of the second-order approximation, measured in both the $L^{2}$ and $L^{\infty}$ norms, are substantially smaller than those of the leading-order approximation. Moreover, the $L^{\infty}$ norm of the correction term $\frac{3}{2}(1-4z^{2})(30z^{2}-\tfrac{3}{2})u_{2}(y)$ is $0.0249$, which is of the same order of magnitude as the error in the leading-order approximation. This indicates that the correction term provides a practical error estimate of the leading-order approximation.
The flux of this flow approximation (26) is
$F$ $=\int\limits_{-1}^{1}\int\limits_{-\frac{L}{2}}^{\frac{L}{2}}\frac{3}{2}\left(1-4z^{2}\right)\left(u_{0}+\left(30z^{2}-\frac{3}{2}\right)u_{2}\right)\mathrm{d}y\mathrm{d}z$ (27)
$=\sum\limits_{j\in\left\{0,2\right\}}a_{0j}a_{0j}\left(L-2\sqrt{\lambda_{j}}\text{tanh}\left(\frac{L}{2\sqrt{\lambda_{j}}}\right)\right)$
$=L-\frac{1}{5}\sqrt{\frac{1}{266}\left(9\sqrt{7}+2618\right)}+\mathcal{O}\left(e^{-\frac{L}{\sqrt{\lambda_{0}}}}\right).$
The constant term in the asymptotic expansion (27) obtained from the second order approximation is $0.63029$, matching the exact flux asymptotic expansion to four digits. In contrast, the leading order approximation matches only to two digits. When $L=1$, the relative error of this flux approximation is $9.47\times 10^{-5}$. For comparison, the relative error of the leading-order flux formula (13) is $6.68\times 10^{-3}$, which is larger by a factor of $70.6$. The error of this second-order flux approximation over a range of aspect ratios $L$ is shown as the orange curve in Figure 2(d), demonstrating uniformly higher accuracy across all tested domain sizes.
Figure 5: Simulation of the centrifuge-on-a-chip. The computational domain is $(x,y,z)\in[-5,5]\times[-5,5]\times[-0.5,0.5]$. The simulation uses a Reynolds number of 100 and P2–P1 finite elements. (a) Heatmaps of the $x$-components of $\mathbf{u}_{0}$ (top) and $\mathbf{u}_{2}$ (bottom) obtained by projecting the 3D field onto the corresponding basis functions. (b) Heatmaps of $w_{2}$ from the 3D simulation (top) and from the 2D reduced model (bottom). (c) Transects of the $x$-velocity near the outlet ($x=4.9$, $y=0.3$): blue dashed curve shows reconstruction using $\mathbf{u}_{0}$ only; orange dot dashed curve shows reconstruction using $\mathbf{u}_{0}$ and $\mathbf{u}_{2}$ via (20); black solid curve shows the full 3D result. (d) Transects of $w$ at x=1x=-1, $y=0.25$ from the 3D simulation (black) and from the 2D reduced model (orange dashed). Vertical velocity units are scaled by $10^{-3}$.
Next, we apply the second-order approximation (23) to the centrifuge-on-a-chip in figure 1 (c). (26) is used as the boundary condition at the inlet in the simulation of the 2D approximate equation. To ensure the flux is 1 at the inlet, we normalize (26) using (27).
Second-order corrections are especially valuable when the leading order approximation suffers large errors, or equivalently, when the flow being approximated is highly non-parabolic, or has a cross-gap component. We analyze the centrifuge-on-a-chip to determine where these sources of error occur. In figure 5 (a), we plot the magnitude of $\mathbf{u}_{0}$ and $\mathbf{u}_{2}$ calculated by projecting 3D simulation to the basis functions for $\mathbf{u}_{0}$ and $\mathbf{u}_{2}$. The relative size of those components gives a measure of how non-parabolic the flow is. In figure 5 (b), top panel, we plot the 3D simulated $w$-field. Taken together, the 3D simulations reveal that the leading-order 2D model becomes inaccurate near the junction between the camber and the outlet. In this region, the gapwise velocity profile departs significantly from the parabolic shape assumed by the leading-order theory, and the cross-gap velocity attains its largest magnitude over the entire domain. These effects indicate the breakdown of the leading-order approximation, and motivate the need for including second-order corrections.
Including the next-order terms alleviates both deficiencies. The bottom panel of figure 5(b), computed from the second-order 2D equations, shows good overall agreement with the full 3D simulation. Figures 5(c) and 5(d) present the in-plane and cross-gap velocity profiles as functions of $z$ at a representative point in the domain, demonstrating close quantitative match between the second-order approximation and the 3D results.
To quantitatively validate the approximation, we compare the produced separation eddy size in figure 4, finding that the extended 2D approximation capture the eddy size, which is useful for predicting separation performance with almost imperceivable error.
@@
What general benefit can we gained from higher-order approximations? First, in many microfluidic thin-gap flows, the vertical velocity remains small, making truly 3D flows rare in thin-gap geometries. Conversely, truly 2D flows occur only in special cases, such as Stokes flow in a channel with constant radius of curvature Lauga et al. (2004). Some applications, however, deliberately induce strong three-dimensionality, e.g. micromixing Stroock et al. (2002). The higher-order approximation (20) provides a nontrivial description of the vertical velocity component, offering valuable insight in these contexts. Second, higher-order terms provide a built-in mechanism for estimating and controlling approximation error. If the estimated error exceeds a chosen tolerance, additional terms can be systematically added to improve accuracy.
@@
While our analysis primarily focused on approximating fluid flows. So long as the domain has a thin dimension, the approach is flexible and broadly applicable. It can be readily extended to other physical quantities, such as solute concentration, electric fields, and magnetic fields, arising in applications like reactive solute transport (Salmon & Ajdari (2007)), electrokinetic flows (Lin et al. (2008); Ding (2023)), and magnetohydrodynamic microfluidics (Qian & Bau (2009)). The method also allows for generalization of boundary conditions, for example, incorporating no-stress boundaries to model bubble motion (Booth et al. (2025)) or flows in microchannels with soft walls (Inamdar et al. (2020)).
5 Acknowledgments
@@
6 Declaration of interests
The authors report no conflict of interest.